OpenStudy (anonymous):

How many molecules are in 4.75 moles H2SO4?

OpenStudy (aaronq):

1 mole = 6.023x10^23 molecules

OpenStudy (anonymous):

I know that, but I'm confused on how to convert it.

OpenStudy (aaronq):

if you have 4.75 dozens of doughnuts, you have 4.75 x 12= if you have 4.75 moles of molecules, you have 4.75 x avogadros number

OpenStudy (anonymous):

also do you know why it ends up as like 10^4? example: 4.84 * 10^4. Why does the power of 10 change?

OpenStudy (aaronq):

it shouldn't go down

OpenStudy (aaronq):

you should get 2.86x10^24

OpenStudy (anonymous):

but 4.75*6.022= 28

OpenStudy (aaronq):

it's not 6.023, it's 6.023x10^23 as in 602300000000000000000000

OpenStudy (anonymous):

can you show me what variables to multiply because I'm still very confused.

OpenStudy (aaronq):

you have 4.75 moles right? so multiply 4.75 x (6.023x10^23)=

OpenStudy (anonymous):

don't we have to find the MM of H2SO4 and multiply it by that first?

OpenStudy (anonymous):

molar mass

OpenStudy (aaronq):

no they're asking you for the amount of molecules not grams

OpenStudy (anonymous):

okay, but does (4.75 x 6.023) multiply with 10^23?

OpenStudy (aaronq):

do you know what scientific notation is? avogadros number is in scientific notation 6.023x10^23, is one number not 2.

OpenStudy (anonymous):

wouldn't 6.023 x 10^23 be some outrageous number

OpenStudy (anonymous):

WAIT

OpenStudy (anonymous):

I think I understand

OpenStudy (aaronq):

yes it is, because molecules are really, like really small

OpenStudy (anonymous):

say we multiply 4.75 x 602300000000000000000000 (something like that) how does it end up with 10^24?

OpenStudy (aaronq):

lol pretend it wasn't 10^23 pretend it was 6023 (six thousand and twenty three) multiply it by 4.75 4.75*6023=28609.25 you went from 6.023x10^3 to 2.8x^10^4 one order of magnitude change same thing with the problem you're doing

OpenStudy (anonymous):

oh alright that makes sense, but if we want to solve this equation again do we have to always multiply by 6023000000000000000000?

OpenStudy (anonymous):

isn't there a simpler way?

OpenStudy (aaronq):

if you're finding the number of atoms, molecules or ions, you have to use avogadros number

OpenStudy (aaronq):

it's simpler if you use scientific notation 6.023*10^23

OpenStudy (anonymous):

I mean you have to times by 6.023 x 10^23, but how do you use the scientific notation?

OpenStudy (aaronq):

that is scientific notation 6.023*10^23, in most calculators you can input the number like that

OpenStudy (anonymous):

Alright, thank you for your help. Will it be alright if I happen to message you again if I have any questions?

OpenStudy (aaronq):

yeah, for sure. i have a crappy internet connection right now, so it might take me a bit, but il reply

OpenStudy (anonymous):

alright thanks again for your help!

OpenStudy (aaronq):

no problem !